# Dynamic Strategy Planning for Efficient Question Answering with Large Language Models

### v1

@@ 8 Conclusion and Future Work
In our work, we introduce the paradigm of dynamic strategy planning for question-answering mimicking human cognitive thinking through DyPlan with the goal of reducing inference costs and improving model performance. By adding verification and self-correction using DyPlan-verify, we further enhance the model output quality. Through experimentation on three MHQA datasets, we show strong efficacy and improved performance using our techniques. Our analyses and empirical bounds provide promise for further improvements. Incorporating partial thinking and integrating dynamic tool usage can be explored to further improve DyPlan. Utilizing alignment-based fine-tuning can further improve the model’s effectiveness.
@@ A.2 Fixed Strategy Implementations
We self-implemented the simple Direct strategy. We utilized the codebase [^4] https://github.com/oneal2000/DRAGIN of DRAGIN Su et al. (2024) to implement the Chain-of-Thought (CoT) Wei et al. (2022) and RAG Lewis et al. (2020) strategies. We utilized a BM25 retrieval system indexed on the entire Wikipedia and capable of retrieving intermediate excerpts of length 200 based on the query. We provide the top three passages as the retrieved passages for RAG. For SelfAsk Press et al. (2023), we utilized their original codebase. [^5] https://github.com/ofirpress/self-ask If any of the strategy inferences weren’t able to provide their answer within the max generation length limit, we used force-decoding with a preset prefix “Final answer:" to get the final answer. Other specific hyperparameters are provided for each strategy in Tables 678, and  9.
@@ Results:
We provided results for this study for the 2WikiMultihopQA dataset in Table 4. Here, we also provide similar comparisons on HotpotQA and Musique datasets in Tables 14 and 15, respectively. Across all the datasets, we can notice the sub-optimal performance of non-fine-tuned LLM runs with DyPlan. The zero-shot model mostly selects retrieval, while the few-shot model selects other strategies, but it’s not well-calibrated. The calibration is poorer for few-shot DyPlanas the model gets heavily influenced by the in-context examples. In conclusion, we demonstrate how base LLMs by default are not calibrated well to utilize DyPlan, underlining the need for fine-tuning LLMs.
@@ C.2 Qualitative Examples for DyPlan-verify
Which band, Letters to Cleo or Screaming Trees, had more members? Final answer: "Letters to Cleo" Letters to Cleo had 4 members: Brian Desveaux, Michael Hubbard, Ken Ray, and Stacy Jones. Screaming Trees had 4 members: Mark Lanegan, Gary Lee Conner, Van Conner, and Barrett Martin. Both bands had the same number of members, 4. Final answer: "same"

### v2

@@ 8 Conclusion and Future Work
In our work, we introduce the paradigm of dynamic strategy planning for question-answering mimicking human cognitive thinking through DyPlan with the goal of reducing inference computational costs and improving model performance. By adding verification and self-correction using DyPlan-verify, we further enhance the model output quality. Through experimentation on three MHQA datasets, we show strong efficacy and improved performance using our techniques. Our analyses and empirical bounds provide promise for further improvements. Incorporating partial thinking and integrating dynamic tool usage can be explored to further improve DyPlan. Utilizing alignment-based fine-tuning can further improve the model’s effectiveness.
@@ A.2 Fixed Strategy Implementations
We self-implemented the simple Direct strategy. We utilized the codebase [^4] https://github.com/oneal2000/DRAGIN of DRAGIN Su et al. (2024) to implement the Chain-of-Thought (CoT) Wei et al. (2022) and RAG Lewis et al. (2020) strategies. We utilized a BM25 retrieval system indexed on the entire Wikipedia and capable of retrieving intermediate excerpts of length 200 based on the query. We provide the top three passages as the retrieved passages for RAG. For SelfAsk Press et al. (2023), we utilized their original codebase. [^5] https://github.com/ofirpress/self-ask If any of the strategy inferences weren’t able to provide their answer within the max generation length limit, we used force-decoding with a preset prefix “Final answer:" to get the final answer. Other specific hyperparameters are provided for each strategy in Tables 91011, and  12.
@@ Results:
We provided results for this study for the 2WikiMultihopQA dataset in Table 6. Here, we also provide similar comparisons on HotpotQA and Musique datasets in Tables 17 and 18, respectively. Across all the datasets, we can notice the sub-optimal performance of non-fine-tuned LLM runs with DyPlan. The zero-shot model mostly selects retrieval, while the few-shot model selects other strategies, but it’s not well-calibrated. The calibration is poorer for few-shot DyPlanas the model gets heavily influenced by the in-context examples. In conclusion, we demonstrate how base LLMs by default are not calibrated well to utilize DyPlan, underlining the need for fine-tuning LLMs.